OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. x+8y=14 =3x+4y=0

OpenStudy (anonymous):

Help Help!!

OpenStudy (anonymous):

have you ever tried to solve it yourself? Where do you have problem?

OpenStudy (radar):

Is the second equation copied correctly?

OpenStudy (anonymous):

system is: x+8y=14 3x+4y=0 ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i did not mean to put the equal sign in front of second equation

OpenStudy (anonymous):

this one i need some help with and yes of coarse i tried on my own. if you do not mind i would greatly appreciate help with this one. Thank you

OpenStudy (anonymous):

x+8y=14 3x+4y=0 --- x=14-8y * --- * 3(14-8y)+4y=0 --- * 42-24y+4y=0 --- * -20y=-42 y=2,1 --- x+8(2,1)=14 x+16,8=14 x=-2,8 x=-2,8 y=2,1

OpenStudy (anonymous):

solve it the same way we helped you in previous questions... from first equation, you know that x = ###### so you can substitute into second equation 3*(#####) +4y=0 now you know what is y and then you plug y value back to first equation to get what is x

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

fontez let him think a little bit...

OpenStudy (anonymous):

:)